. Z is a standard normal random variable. The P (-1.20 "greater than or equal to" Z "less than or equal to" 1.50) equals _____________. Probability notation for this example P (-1.20 ≤ Z ≥ 1.50) = _________.
Answer #1

Solution :

Given that,  

Using standard normal table ,

=1 - P(-1.20 z 1.50)

= 1 - P(z 1.50) - P(z -1.20)

= 1 - P(0.9332 - 0.1151)

= 1 - 0.8181

= 0.1819

P(-1.20 < z < 1.50) = 0.1819